MFT(1) MFT(1) NAME mft−translateMetafontcodetoTeXcodeforprettyprinting SYNOPSIS mftmffilename[.mf][-cs][changefilename[.ch]][stylefilename[.mft]] DESCRIPTION ThemftprogramcreatesaTeXfilefromaMetafontprogram.Ittakesappropriatecareoftypographicdetailslikepagelayoutandtheuseofindentation,italics,boldface,etc.,asillustratedin thebookComputerModernTypefaces,byDonaldKnuth.SpecialconventionsinMetafontcommentsallowyoutocontrolthingsthatwouldnototherwisecomeoutright;section1oftheMFT sourceprogramintheMETAFONTwarereportexplainstheserules. Thecommandlinehasonerequiredfilenameandtwooptionalfilenames.Therequiredoneisa Metafontsourcefile;thereisalsoanoptionalchangefile(whichworksjustasthechangefilesto tangle(1)andweave(1)do)andanoptionalstylefile(whichisprependedtoeverything). Afilenamethatdoesn’tcontainadotisalwaysgivenanextension,either.mf(Metafont)or.ch (change)or.mft(style).Ifnostylefileisspecified,thestylefileplain.mftisautomaticallyused. Thechangefilenameisnotsearchedforusinganypaths.The.mffileissearchedforusingthe MFINPUTSenvironmentvariableifyouhavesetit,orelsethesystemdefault: .:/usr/TeX/lib/mf/macros:/usr/TeX/src/TeX+MF/typefaces//.Thestylefileis searchedforusingtheTEXINPUTSenvironmentvariable;thedefaultthereis: .:/usr/TeX/lib/tex/macros//.Seetex(1)forthedetailsofthesearching. TheoutputTeXfilenameisformedbyusing.texinplaceoftheextensionofmffilename. FILES /usr/TeX/lib/tex/macros/mftmac.texTeXmacrosusedbymftoutput. /usr/TeX/lib/tex/macros/plain.mftDefaultstylefile. /usr/TeX/lib/tex/macros/cmbase.mftStylefileforComputerModern. SEEALSO mf(1),weave(1) DonaldKnuthetal.,METAFONTware AUTHORS DonaldKnuthwrotetheprogram,andheportedittoUnixwiththehelpofPierreMacKayand theUnixportofweavebyHowardTrickeyandPavelCurtis.Theprogramispublishedinthe METAFONTwaretechnicalreport,availablefromtheTeXUsersGroup. 7Jan92 1